OpenStudy (anonymous):

if you get an average of 145 and you want to get the average to 150 how would you get it there (what number would you add to it)

OpenStudy (anonymous):

i added the numbers...158,124,110,167,146,and165 and got 870 then i divided it by 6 and got 145 now what other number do i have to add to get to 150

OpenStudy (anonymous):

If want to get 150, using two numbers your answer is 155. However with more numbers you use the same method, except divide by however many numbers you now have

OpenStudy (amistre64):

lets say your original amount is equal to A such that: A/6 = 145 you want (A+x)/7 = 150 150(7) - A = x

OpenStudy (anonymous):

what is the (7) for what do yyou do with it

OpenStudy (amistre64):

since your adding an amount; your number of items increase from 6 to 7 right?

OpenStudy (amistre64):

i see some short cut tho .... A/6 = 145 A = 145(6) 150(7) - 145(6) = x

OpenStudy (amistre64):

so im going with 180 :)

OpenStudy (amistre64):

now, what of part of that makes sense and what doesnt?

OpenStudy (anonymous):

\[\frac{145+x}{n}= 150\]\[x=150 n-145 \]

OpenStudy (anonymous):

n is the number of addends used to calculate the original total. Let n = 5. Then 145*5, 725 is the original total. 150*5-145=605 (145+605)/5 = 150